OpenStudy (anonymous):

Find the 22nd term: a1 = 6; d = 4

OpenStudy (anonymous):

90 is the answer ryt ?????

OpenStudy (anonymous):

yep, ty :)

OpenStudy (anonymous):

do u wan the steps ????

OpenStudy (anonymous):

yes please :)

OpenStudy (anonymous):

a22=a+(n-1)d = a+21d =6+21*4 =6+84 =90

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):